Problem 14-6 Calculating Cost of Debt [LO2]

Viserion, Inc., is trying to determine its cost of debt. The firm has a debt issue outstanding with 13 years to maturity that is quoted at 104 percent of face value. The issue makes semiannual payments and has an embedded cost of 5 percent annually.

a.

What is the company’s pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. If the tax rate is 22 percent, what is the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
a. Pretax cost of debt %
b. Aftertax cost of debt %
Answer #1

Ans:- (a) In this part we need to find the pretax cost of Debt.

In this question, the Pretax Cost of debt is nothing but the yield to maturity of the bond.

Par value is given $1000 and Bond Price will be $1000 * 104% =$1040.

Now the semi-annual coupon value of the bond will be (1000 * 5%)/2 = 25.

It is given 13 years to maturity that means the maturity period will be 13 * 2 = 26.

Now the Yield To Maturity can be easily calculated by the Rate function of excel.

=RATE(nper,pmt,pv,fv)

=RATE(26,25,-1040,1000)

=2.294%.

The semi-annual yield to maturity is 2.294%. Now the annual yield to maturity will be 2.294*2 = 4.59%.

Therefore The Companys Pretax Cost of the debt is 4.59%.

(b) Now the after-tax cost of debt will be given by Pretax cost of debt * ( 1 - Tax rate). The tax rate is given 22%.

Therefore the After-tax cost of debt will be 4.59 * ( 1 - 0.22)

= 3.5802%.

Note:- Please give thumbs up if this answer helps you thankyou very much.